OpenStudy (anonymous):

The areas of two similar rectangles are 338 cm^2 and 72 cm^2. What is the similarity ratio of the perimeters? A. 338:72 B. 169:36 C. 29:12 D. 13:6

OpenStudy (nurali):

338/72 = 169/36 sqrt(169)/sqrt(36) = 13/6